OpenStudy (anonymous):

Factor the following completely Im not sure how to do this please show work I will give medals 1. c^2 – 121 2. p^2 + 22p + 40

OpenStudy (whpalmer4):

I told you how to do both of these when you posted them before. Please do not repost if you aren't going to heed what you are told.

OpenStudy (anonymous):

sorry i was still confused

OpenStudy (whpalmer4):

Then ask questions! Here's the first one: \[a^2-b^2 = (a-b)(a+b)\] Compare \[a^2-b^2\]\[c^2-121\]what is the square root of 121? Are those expressions not equivalent if \(a=c\) and \(b = \sqrt{121}\)?

OpenStudy (whpalmer4):

For the second one, use the splitting or grouping technique I showed you earlier: \[p^2 + 22p + 40\]Multiply the coefficient of the first and last terms: 1*40 = 40 Now choose a pair of factors of 40 that add to 22: 2*20 = 40, 2+20 = 22 Rewrite the polynomial, splitting the middle term:\[p^2 + 20p + 2p + 40\]Now group each pair\[(p^2+20p) + (2p+40)\]Now factor each group; what do you get?

OpenStudy (anonymous):

so the first ones answer would be (c - 11)(c + 11)?

OpenStudy (whpalmer4):

Yes!

OpenStudy (whpalmer4):

Whenever you see <something> - <something else>, you should always be investigating whether both things are squares like this. It's just about the easiest factoring of all.

OpenStudy (anonymous):

and the second ones answer would be (p + 20)(p + 2)?

OpenStudy (whpalmer4):

Yes again! \[(p^2+20p) + (2p+40)\]\[p(p+20) + 2(p+20)\]\[(p+20)(p+2)\] Checking our work: \[(p+20)(p+2) = p^2 + 2p + 20p + 20*2 = p^2 + 22p + 40\checkmark\]

OpenStudy (whpalmer4):

Checking the other one, for good measure: \[(c+11)(c-11) = c^2 -11c + 11c -121 = c^2 -121\checkmark\]

OpenStudy (anonymous):

thank you

OpenStudy (whpalmer4):

You're welcome. I hope you're now a factoring master :-)